# Is 1:1 Always Most Powerful? Why Unequal Allocation Merits Broader Consideration

### v1

@@ 4 Discussion
Whether and, crucially, by how much unequal allocation can increase statistical power for a given statistical test depends on several key factors:
@@
While Neyman allocation might indicate a departure from a 1:1 ratio is optimal, the benefits in terms of statistical power become less pronounced when a study is already robustly powered with ER. This may be the case for studies with a large sample size or an expected large treatment effect. Our key objective, however, was to highlight that implementing unequal allocation is not inherently detrimental to statistical power while it may be highly desirable for other reasons. A reduction in power (compared to ER) would only occur if the chosen allocation ratio diverged from the direction Neyman allocation would prescribe or exceeding it too far, e.g. in Case Study 2 choosing a 2:1 ratio ($76.2\%$ power) or 1:19 ratio ($87.3\%$ power).
Pinpointing the optimal unequal ratio for FUR at the trial’s outset is challenging. This hurdle can be overcome by using RAR instead. The use of RAR to target Neyman allocation in Case Study 1 leads to a more substantial benefit, boosting power by $10$ percentage points. However, using this type of RAR for binary endpoints in combination with the Wald test can lead to serious type-I error inflation. Methods exist to control type-I error when using RAR (Pin et al., 2025). In this case study the RAR design incorporates a “burn-in" period where the first 12 patients are allocated equally, after which the randomization probabilities are adapted. For Case Study 2 with the continuous endpoint and relatively large sample size RAR targeting Neyman does control type-I error and almost reaches the optimal allocation ratio of 1:4 ($n_{1}/n=0.798$ instead of $0.8$). We used a burn-in of 70 patients in total.
While power increased in our examples, FUR or RAR aiming for Neyman allocation can also yield power comparable to ER without significant gains, especially when the optimal allocation is near 1:1. For binary endpoints, Figure 1 shows optimal Neyman allocation deviates little from equal allocation when success probabilities are close to 0.5. Unequal allocation offers the most advantage in terms of statistical power when success probabilities are dissimilar and not close to around 0.5.
It is important to note that our findings are based on a two-arm setting using the parametric Wald test. Choosing an optimal allocation ratio for more specific or complex scenarios and other tests requires further study, (e.g. Jackson and Cox, 2024).
Ultimately, the potential benefit of unequal allocation should be assessed on a case-by-case basis for any given trial design and statistical test. Our analysis demonstrates that considering the Neyman allocation for a study can bring advantages beyond improving statistical power; these gains can translate into significant economic benefits and, most importantly, improved patient outcomes. Consequently, we recommend that trial designs move beyond a default 1:1 ratio. Unequal allocation warrants consideration, as it may offer benefits even when statistical power is on par with a 1:1 approach.

### v2

@@ 4 Discussion
Whether and, crucially, by how much equal allocation can be superior in terms statistical power for a given statistical test depends on several key factors:
@@
While Neyman allocation might indicate a departure from a 1:1 ratio is optimal, the benefits in terms of statistical power become less pronounced when a study is already robustly powered with ER. This may be the case for studies with a large sample size or an expected large treatment effect. Our key objective was to demonstrate that implementing an unequal allocation ratio—often highly desirable for patient benefit —is not inherently detrimental to statistical power. We found that a substantial reduction in power, relative to equal randomization, occurs only when the chosen allocation ratio deviates markedly from Neyman allocation, and not in an already overpowered study. In particular, power losses are observed only when a sub-optimal ratio is skewed in the wrong direction or excessively unbalanced—for example, a 2:1 allocation (yielding 85.4% power) or a 1:9 allocation (80.8% power) in our first case study—compared with 91.7% power under equal randomization (see Table 1).
We found that a substantial reduction in power (relative to equal randomization) occurs only when the chosen allocation ratio substantially misses the optimal (Neyman) proportion and not in an overpowered study. Specifically, choosing a sub-optimal ratio that is skewed in the wrong direction or excessively skewed, such as a 2:1 ratio (resulting in $85.4\%$ power) or 1:9 ratio ($80.8\%$ power) in our first case study, is required to see significant penalties compared to 91.7% power under ER, see Table 1.
While power increased in our examples, FUR or RAR aiming for Neyman allocation often yields power comparable to ER, especially when the optimal allocation is near 1:1. For binary endpoints, Figure 1 shows that optimal Neyman allocation deviates little from equal allocation when success probabilities are close to 0.5. In many regions of the parameter space, power gains are minimal or nonexistent, but it can still favour participants by assigning more to the better-performing treatment with minimal impact on power. Thus, the value of unequal allocation lies not just in maximizing power, but also in enhancing patient outcomes without decreasing statistical power. It is important to note that our findings are based on a two-arm setting using the parametric Wald test. Choosing an optimal allocation ratio for more specific or complex scenarios and other tests requires further study, (e.g. Jackson and Cox, 2025).
Beyond statistical power and participant benefit, unequal allocation may also offer financial and practical advantages. First, if an unequal ratio reduces the total sample size needed to achieve a given power, this directly lowers trial costs. Second, when treatment arms differ in per-patient costs, favouring the less expensive arm can further reduce overall expenditure. From a patient perspective, allocating more participants to the better-performing treatment can reduce burden and, more generally, make trial participation more appealing, potentially speeding enrolment (Meurer et al., 2012; Tehranisa and Meurer, 2014). Practically, unequal ratios can also help accommodate real-world constraints such as limited drug supply or differential treatment availability across centres. These factors provide compelling advantages for unequal allocation, and these benefits are often ignored due to the misconception that the practice is inherently detrimental to statistical power.
The potential benefit of unequal allocation extends well beyond simple statistical power. Our analysis demonstrates that failing to consider unequal allocation (e.g. Neyman allocation) —even when its power advantage is minimal— often means sacrificing significant gains in patient welfare and potential economic efficiency. Consequently, we advocate for a fundamental shift in practice: trial designs must move beyond the default 1:1 ratio toward the deliberate determination of the optimal allocation ratio based on trial-specific parameters. The allocation ratio should always be an explicit, data-driven decision assessed against the full spectrum of statistical, ethical, and financial benefits available to the study.
@@ Acknowledgements
  • We thank Alex John London and Stephen Senn for their insightful comments on an earlier draft of this paper.